In a round-robin tennis tournament, every player meets every other player exactly once. How many players can participate in a tournament of 45 matches?
Answer by scott8148(6628)

let x="number of players" __ each player plays x-1 matches BUT a vs b is the same as b vs a
__ so the number of matches is x(x-1)/2

x(x-1)/2=45 __ multiplying by 2 and subtracting 90 __ x^2-x-90=0

factoring __ (x+9)(x-10)=0 __ x+9=0 and x-10=0